kierrxn:

Which represents the solution(s) of the system of equations, \(y = x^2 – 6x + 8\) and \(y = –x + 4\)? Determine the solution set by graphing. A. (1, 3) only B. (2, 0) and (4, 0) C. (1, 3) and (4, 0) D. No solutions

Enorym:

You can set them equal to each other (Substitution since they are both equal to y) \( x^2 - 6x + 8 = -x + 4\) You can set this all equal to zero (I will assume you know how to do this) \(x^2 - 5x + 4 = 0\) From here you can solve by factoring or completing the square If you have any questions feel free to ask (: